How do you solve the following inequality?
2+3(m+5)≥ 4(m+3)
(the answer key says (m|m ≤ 5 ))
i kept getting -13 ≥ m ?
thank you!

Found 2 solutions by Theo, lwsshak3:
Answer by Theo(13342)

your equation is:
2+3(m+5)≥ 4(m+3)
simplify by multiplying out what's in the parentheses to get:
2 + 3m + 15 >= 4m + 12
subtract 4m from both sides of the equation to get:
2 - m + 15 >= 12
subtract 15 from both sides of the equation to get:
2 - m >= -3
subtract 2 from both sides of the equation to get:
-m >= -5
multiply both sides of the equation by -1 to get:
m <= 5
this solution can also be shown as:
{set of all m such that m is an element of the real numbers and m <= 5}
that's equivalent to what you show as:
(the answer key says (m|m ≤ 5 ))

Answer by lwsshak3(11628)

How do you solve the following inequality?
2+3(m+5)≥ 4(m+3)
2+3m+15≥4m+12
3m-4m≥-17+12
-m≥-5
divide by -1 and reverse inequality sign
m≤5
solution: {m|m≤5}
or (-∞,5] (interval notation)